# Positive Semidefinite Supermartingales and Randomized Matrix Concentration Inequalities

### v3

@@ C.1 Doob’s $L^{p}$ Inequality on Trace
Fact 3 and Fact 4, on the other hand, convert a matrix-valued (super/sub)martingale to a scalar-valued one. For example,
@@
Let $\{\mathbf{M}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued submartingale or $\mathcal{S}_{d}$-valued martingale, and $p\geqslant 1$. Then, $\{\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n})^{p}\}$ is a $\mathbb{R}_{\geqslant 0}$-valued submartingale.
@@
If $\{\mathbf{M}_{n}\}$ is a martingale, the right hand side above is $\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n-1})^{p}$. If $\mathbf{M}_{n}\succeq 0$ is a submartingale, then we have by Fact 3,
@@
Proposition C.3 (Trace Doob’s $L^{p}$ Inequality).
Let $\{\mathbf{M}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued submartingale or $\mathcal{S}_{d}$-valued martingale, and $p\geqslant 1$. Then,
$\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}\|\mathbf{M}_{n}\|_{\mathrm{op}}^{p}\right)\leqslant\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}(\mathrm{tr}\operatorname{abs}\mathbf{M}_{n})^{p}\right)\leqslant\left(\frac{p}{p-1}\right)^{p}\mathbb{E}(\mathrm{tr}\operatorname{abs}\mathbf{M}_{N})^{p}\leqslant\left(\frac{pd}{p-1}\right)^{p}\mathbb{E}\|\mathbf{M}_{N}\|_{\mathrm{op}}^{p}.$ (157)
@@
The second inequality is directly from applying Doob’s $L^{p}$ inequality on the nonnegative submartingale $\{\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n})\}$. The first and the third $\leqslant$’s are both trivial. ∎
Appendix D Non-symmetric, Non-square Matrices
@@
In Table 2, we see that quadratic assumptions on the possibly non-square random matrix $\mathbf{X}_{1}$ translate directly to those on the symmetric random matrix $D(\mathbf{X}_{1})$.
@@
$\mathbf{X}_{1}^{\mathsf{T}}\mathbf{X}_{1}\preceq\mathbf{B}_{1}$ and $\mathbf{X}_{1}\mathbf{X}_{1}^{\mathsf{T}}\preceq\mathbf{B}_{2}$ $D^{2}(\mathbf{X}_{1})\preceq\left(\begin{matrix}\mathbf{B}_{1}&\\ &\mathbf{B}_{2}\end{matrix}\right)$
@@
$\mathbf{Y}_{n}=\mathrm{e}^{\mathbf{F}_{1}/2}\mathrm{e}^{\mathbf{E}_{1}/2}\mathrm{e}^{\mathbf{F}_{2}/2}\mathrm{e}^{\mathbf{E}_{2}/2}\dots\mathrm{e}^{\mathbf{F}_{n}/2}\mathrm{e}^{\mathbf{E}_{n}/2}\mathrm{e}^{\mathbf{E}_{n}/2}\mathrm{e}^{\mathbf{F}_{n}/2}\dots\mathrm{e}^{\mathbf{E}_{2}/2}\mathrm{e}^{\mathbf{F}_{2}/2}\mathrm{e}^{\mathbf{E}_{1}/2}\mathrm{e}^{\mathbf{F}_{1}/2}.$ (160)
is a $\mathcal{S}_{d_{1}+d_{2}}^{+}$-valued supermartingale.

### v4

@@ C.2 A Matrix Doob’s $L^{p}$ Inequality
Fact 2 and Fact 3, on the other hand, convert a matrix-valued (super/sub)martingale to a scalar-valued one by trace scalarization. For example, denoting by $\|\mathbf{X}\|_{p}$ the Schatten $p$-norm of $\mathbf{X}\in\mathcal{S}_{d}$; that is, $\|\mathbf{X}\|_{p}=(\mathrm{tr}((\operatorname{abs}\mathbf{X})^{p}))^{1/p}$.
@@
Let $\{\mathbf{M}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued submartingale or $\mathcal{S}_{d}$-valued martingale, and $p\geqslant 1$. Then, $\{\|\mathbf{M}_{n}\|^{p}_{p}\}$ is a submartingale.
@@
If $\{\mathbf{M}_{n}\}$ is a martingale, the right hand side above is $\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n-1})^{p}$. If $\mathbf{M}_{n}\succeq 0$ is a submartingale, then we have by Fact 2,
@@
In particular, when $p=1$ in above, the nuclear norm process $\{\|\mathbf{M}_{n}\|_{1}\}$ is a submartingale. We now have the following generalization of Doob’s $L^{p}$ maximal inequality (e.g. Theorem 12.1 in Klenke [2013]).
Proposition C.3 (Doob-type Nuclear $L^{p}$ Inequality).
Let $\{\mathbf{M}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued submartingale or $\mathcal{S}_{d}$-valued martingale, and $p>1$. Then,
$\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}\|\mathbf{M}_{n}\|^{p}\right)\leqslant\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}\|\mathbf{M}_{n}\|_{1}^{p}\right)\leqslant\left(\frac{p}{p-1}\right)^{p}\mathbb{E}\|\mathbf{M}_{N}\|^{p}_{1}=\left(\frac{p}{p-1}\right)^{p}\mathbb{E}(\|\mathbf{M}_{N}\|\cdot\operatorname{r}(\mathbf{M}_{N}))^{p}.$ (140)
@@
The second inequality is directly from applying Doob’s $L^{p}$ inequality on the nonnegative submartingale $\{\|\mathbf{M}_{n}\|_{1}\}$. The first $\leqslant$ and the final $=$ follow from straightward relations between the spectral and nuclear norms, and the effective rank. ∎
C.3 Non-symmetric, Non-square Matrices
@@
In Table 2, we see that quadratic assumptions on the possibly non-square random matrix $\mathbf{X}_{1}$ translate directly to those on the symmetric random matrix $D(\mathbf{X}_{1})$. Conditional distribution assumptions are translated similarly.
@@
$\mathbf{X}_{1}^{\mathsf{T}}\mathbf{X}_{1}\preceq\mathbf{B}_{1}$ and $\mathbf{X}_{1}\mathbf{X}_{1}^{\mathsf{T}}\preceq\mathbf{B}_{2}$ $D^{2}(\mathbf{X}_{1})\preceq\left(\begin{matrix}\mathbf{B}_{1}&\\ &\mathbf{B}_{2}\end{matrix}\right)$ and thus
$\left(\begin{matrix}-\mathbf{B}_{1}^{1/2}&\\ &-\mathbf{B}_{2}^{1/2}\end{matrix}\right)\preceq D(\mathbf{X}_{1})\preceq\left(\begin{matrix}\mathbf{B}_{1}^{1/2}&\\ &\mathbf{B}_{2}^{1/2}\end{matrix}\right)$
$\|\mathbf{X}_{1}\|\leqslant b$ $\|D(\mathbf{X}_{1})\|\leqslant b$
@@
$\mathbf{Y}_{n}=\mathrm{e}^{\mathbf{F}_{1}/2}\mathrm{e}^{\mathbf{E}_{1}/2}\mathrm{e}^{\mathbf{F}_{2}/2}\mathrm{e}^{\mathbf{E}_{2}/2}\dots\mathrm{e}^{\mathbf{F}_{n}/2}\mathrm{e}^{\mathbf{E}_{n}/2}\mathrm{e}^{\mathbf{E}_{n}/2}\mathrm{e}^{\mathbf{F}_{n}/2}\dots\mathrm{e}^{\mathbf{E}_{2}/2}\mathrm{e}^{\mathbf{F}_{2}/2}\mathrm{e}^{\mathbf{E}_{1}/2}\mathrm{e}^{\mathbf{F}_{1}/2}$ (143)
is an $\mathcal{S}_{d_{1}+d_{2}}^{+}$-valued supermartingale.